# Revisiting thermodynamic topology of Hawking-Page and Davies type phase transitions

### v1

@@ Abstract
In the early 1970s, the groundwork in black hole thermodynamics established the basis for comprehending the correlation between black holes and the laws of thermodynamics [1, 2, 3, 4]. Subsequent research endeavors have unveiled numerous compelling findings in this field [5, 6, 7, 8, 9, 10, 11]. One very interesting aspect in the study of black hole thermodynamics is the issue phase transition[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Black holes undergo various types of phase transitions, including Davies type phase transition[12], Hawking–Page phase transition [13], extremal phase transition ( transition of black holes from non-extremal to extremal states) [14, 15, 16, 17, 18, 19, 20, 21, 22] and van der Waals-type phase transition (a phase transition reminiscent of van der Waals behavior)[23, 24, 25, 26, 27, 28, 29, 30]. Phase transitions of black holes have been extensively studied using different frameworks such as extended black hole thermodynamics [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], restricted phase space thermodynamics and holographic thermodynamics [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], geometrical interpretation of phase transition [55, 56, 57, 58, 59, 60, 61, 62, 63, 64] etc. A relatively recent development in this direction is the topological interpretation of black hole phase transition [66, 65, 71, 72, 73, 74, 75, 67, 76, 77, 78, 79, 70, 80, 81, 82, 83, 84, 85, 86, 69, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 68] in which the critical points are endowed with some topological charges. Some of these works have used thermodynamic topology to understand the van der Waals-type phase transition [65, 71, 72, 73, 74] , the Hawking-Page phase transition [67, 76, 77] and Davies-type phase [68]. These existing literatures suggest that topologically investigating Hawking-Page and Davies-type phase transition typically requires defining two separate potentials or vector fields [69, 67, 68, 70]. The key inspiration behind thermodynamic topology of black holes lies in Duan’s $\phi$-mapping current theory[97, 98]. A simple and brief discussion of this approach is provided below:
@@
This suggest that $j^{\mu}$ is non-zero only at the point where $\phi^{a}(x^{i})=0$ where, the $i$-th solution is denoted as $\vec{x}=\vec{z}_{i}$. Using the $\delta$-function theory [Schouten1951], the density of topological current can be obtained as
@@ III Schwarzchild Ads Black hole
Using eq.12 and eq.13, We calculate the free energy as :
@@
Figure 1: Figure (a) represents $F$ vs $r_{+}$ plot for Schwarzchild AdS black hole. The blue dot represents $r_{+}=1$, the free energy becomes zero. This is the exact point at which the Hawking-Page transition takes place. Figure (b) represents $C$ vs $r_{+}$ plot. The black dot represents at $r_{+}=\frac{1}{\sqrt{3}}$, the specific heat diverges. This is the exact point at which Davies-type transition takes place
The exact point at which the Hawking-Page transition takes place is found by setting $F=0$. As shown in the Figure. 1(a), at $r_{+}=1$, Hwking-Page transition takes place.
@@
at $S=\frac{1}{\sqrt{3}}$ the specific heat diverges. This is the exact point at which Davies type transition takes place as shown in figure.1(b).
Now applying eq.14 in eq. 8, we get the component as :
@@
We plot a portion of the vector field ($n^{1},n^{2}$) in Figure.2. As we can see from Figure.2, there are two critical points or zero points in that particular portion of the vector field ($n^{1},n^{2}$). One is at $Z_{D}=0.57735$ and the other at $Z_{H}=1$. Comparing with Figure.1, we can see $Z_{D}$(enclosed by the black contour) is the Davies point at which Davies type phase transition occurs and $Z_{H}$ is the point(enclosed by the blue contour) at which Hawking-Page phase transition occurs.
@@
To calculate the topological charge, we perform a contour integral around the black and the blue colored contour. In Figure.3(a), as represented by the black solid line, the topological charge corresponds to Davis point is found to be $-1$ and the topological charge for Hawking-Page phase transition is $1$ as represented by the blue solid line in Figure.3(b).
Figure 3: Schwarzchild AdS black hole: Figure (a) represents topological number calculation for Davies point around the black contour in Figure.2 and Figure (b) represents the topological number calculation for Hawking-Page phase transition point around the blue contour in Figure.2
@@ IV RN AdS black hole in grand canonical ensemble
The Smarr formula for RN black hole is given as
$m=\frac{q^{2}+r_{+}^{4}+r_{+}^{2}}{2r_{+}}$ (19)
where, we have scaled charge $Q$, and mass $M$ as
$q=Q/l^{2},\quad\text{and}\quad m=M.$ (20)
@@
$q=r_{+}\phi.$ (21)
Also, the mass is given as
@@
The temperature can be calculated from the mass as
@@
The free energy of RN AdS black hole in the grand canonical ensemble is given as
$F=M-ts=-\frac{\sqrt{s}\left(s+\pi\left(\phi^{2}-1\right)\right)}{4\pi^{3/2}}.$ (24)
The plot of free energy is shown in 4(a). The figure shows that the free energy becomes zero at $s=3.11018$ for $\phi=0.1$. This point is identified as the Hawking-Page transition point The specific heat at constant chemical potential is given as
@@
The specific heat for $\phi=0.1$ is shown in 4(b). The black dot represents the Davies point where the specific heat become zero and it is located at $r_{+}=0.57445$. The specific heat diverges for $\phi=\sqrt{1-3r_{+}^{2}}$ which is shown in 4(c). Each point on the curve represents Davies point.
@@
and The normalized vector is represented in Figure.5 for $\phi=0.1$.
@@
The critical points are calculated by setting $\theta=\pi/2$ in $n^{1}$ and equating it to zero. We find two real and positive critical (or zero) points located at $(r_{+},\theta)=(0.57445,\pi/2),(0.99498,\pi/2)$ which are respectively named as $Z_{D}$ and $Z_{H}$. These points exactly match with the Davies points and the Hawking-Page phase transition point. We calculate the topological charge associated with each critical point by encircling contours around them as shown in Figure 6. For the critical point $Z_{D}$ (which corresponds to the Davies point) the topological charge is $-1$ (Figure.6(a))and for the critical point $Z_{H}$ (which corresponds to the Hawking-Page phase transition point) the topological charge is $+1$(Figure.6(a)).
Figure 6: RN AdS black hole : Figure (a) represents topological number calculation for Davies point around the black contour in Figure.5 and Figure (b) represents the topological number calculation for Hawking-Page phase transition point around the blue contour in Figure.5. We have considered $\phi=0.1$.
@@ V Kerr Ads Black hole
The Smarr formula of Kerr Ads black hole in terms of entropy $S$ is given by
@@
Figure 7: Kerr AdS black hole: Figure (a) represents $F$ vs $S$ plot for $\Omega=0.5$. The blue dot represents $S=3.62759$, the free energy becomes zero. This is the exact point at which the Hawking-Page transition takes place. Figure (b) represents $C$ vs $S$ plot for $\Omega=0.5$. The black dot represents at $S=1.20764$, the specific heat diverges. This is the exact point at which the Davies-type transition takes place. In Figure (b) spinodal curve is plotted for different values of $\Omega$.
@@
Here also the topological charge corresponding to the Davies point is determined to be $-1$, while for the Hawking-Page phase transition, it is $1$, as illustrated by the black and blue solid line respectively in Figure.9.

### v2

@@ Abstract
In the early 1970s, the groundwork in black hole thermodynamics established the basis for comprehending the correlation between black holes and the laws of thermodynamics [1, 2, 3, 4]. Subsequent research endeavors have unveiled numerous compelling findings in this field [5, 6, 7, 8, 9, 10, 11]. One very interesting aspect in the study of black hole thermodynamics is the issue phase transition[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Black holes undergo various types of phase transitions, including Davies type phase transition[12], Hawking–Page phase transition [13], extremal phase transition ( transition of black holes from non-extremal to extremal states) [14, 15, 16, 17, 18, 19, 20, 21, 22] and van der Waals-type phase transition (a phase transition reminiscent of van der Waals behavior)[23, 24, 25, 26, 27, 28, 29, 30]. Phase transitions of black holes have been extensively studied using different frameworks such as extended black hole thermodynamics [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], restricted phase space thermodynamics and holographic thermodynamics [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], geometrical interpretation of phase transition [55, 56, 57, 58, 59, 60, 61, 62, 63, 64] etc. A relatively recent development in this direction is the topological interpretation of black hole phase transition [66, 65, 71, 72, 73, 74, 75, 67, 76, 77, 79, 80, 70, 81, 82, 83, 84, 85, 86, 87, 69, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 68] in which the critical points are endowed with some topological charges. Some of these works have used thermodynamic topology to understand the van der Waals-type phase transition [65, 71, 72, 73, 74] , the Hawking-Page phase transition [67, 76, 77, 78] and Davies-type phase [68]. These existing literatures suggest that topologically investigating Hawking-Page and Davies-type phase transition typically requires defining two separate potentials or vector fields [69, 67, 68, 70]. The key inspiration behind thermodynamic topology of black holes lies in Duan’s $\phi$-mapping current theory[98, 99]. A simple and brief discussion of this approach is provided below:
@@
This suggest that $j^{\mu}$ is non-zero only at the point where $\phi^{a}(x^{i})=0$ where, the $i$-th solution is denoted as $\vec{x}=\vec{z}_{i}$. Using the $\delta$-function theory [100], the density of topological current can be obtained as
@@ III Schwarzchild Ads Black hole
Using equations eq.(12) and eq.(13), We calculate the free energy as :
@@
Figure 1: Figure (a) represents $F$ vs $r_{+}$ plot for Schwarzchild AdS black hole. The blue dot represents $r_{+}=1$, the free energy becomes zero. This is the exact point at which the Hawking-Page transition takes place. FIG.(b) represents $C$ vs $r_{+}$ plot. The black dot represents at $r_{+}=\frac{1}{\sqrt{3}}$, the specific heat diverges. This is the exact point at which Davies-type transition takes place
The exact point at which the Hawking-Page transition takes place is found by setting $F=0$. As shown in the FiIG.(1(a),)1 at $r_{+}=1$, Hwking-Page transition takes place.
@@
at $S=\frac{1}{\sqrt{3}}$ the specific heat diverges. This is the exact point at which Davies type transition takes place as shown in FIG.(1(b)).
Now applying equations eq.(14) in eq. (8), we get the component as :
@@
We plot a portion of the vector field ($n^{1},n^{2}$) in FIG.(2). As we can see from FIG.(2), there are two critical points or zero points in that particular portion of the vector field ($n^{1},n^{2}$). One is at $Z_{D}=0.57735$ and the other at $Z_{H}=1$. Comparing with FIG.(1),, we can see $Z_{D}$(enclosed by the black contour) is the Davies point at which Davies type phase transition occurs and $Z_{H}$ is the point(enclosed by the blue contour) at which Hawking-Page phase transition occurs.
@@
To calculate the topological charge, we perform a contour integral around the black and the blue colored contour. In FIG.(3(a),) as represented by the black solid line, the topological charge corresponds to Davis point is found to be $-1$ and the topological charge for Hawking-Page phase transition is $1$ as represented by the blue solid line in FIG.(3(b)).
Figure 3: Schwarzchild AdS black hole: FIG.(a) represents topological number calculation for Davies point around the black contour in FIG.(2), and FIG.(b) represents the topological number calculation for Hawking-Page phase transition point around the blue contour in FIG.(2),
@@ IV RN AdS black hole in grand canonical ensemble
The equation of motion of the RN-AdS black hole in 4-dimension is :
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2})$
where the metric function $f(r)$ is given by :
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{l^{2}}$ (19)
here $M$ and $Q$ is the mass and charge of the black hole respectively and $l$ is the Ads radius. At the event horizon radius $r=r_{+}$ the metric function is set to zero and the mass is obtained as
$M=\frac{l^{2}Q^{2}+l^{2}r_{+}^{2}+r_{+}^{4}}{2l^{2}r_{+}}$ (20)
@@
$\phi=\frac{\partial M}{\partial Q}=\frac{q}{r_{+}}$ (21)
or
$Q=r_{+}\phi$ (22)
Substituting the value of $Q$ in the expression of mass in eq.(20,) we get the new mass in grand canonical ensemble as
@@
where we have set $l=1$ for simplicity. The temperature and the entropy of the black hole is calculated as :
@@
and
$S=\pi r_{+}^{2}$ (25)
The free energy of RN AdS black hole in the grand canonical ensemble is :
$F=M-TS=-\frac{1}{4}r_{+}\left(r_{+}^{2}+\phi^{2}-1\right)$ (26)
The plot of free energy is shown in FIG.(4(a).) The figure shows that the free energy becomes zero at $r_{+}=0.99498$ for $\phi=0.1$. This point is identified as the Hawking-Page transition point he specific heat at constant chemical potential is obtained as :
@@
The specific heat for $\phi=0.1$ is shown in FIG.(4(b)). The black dot represents the Davies point where the specific heat become zero and it is located at $r_{+}=0.57445$. The specific heat diverges for $\phi=\sqrt{1-3r_{+}^{2}}$ which is shown in FIG.(4(c)). Each point on the curve represents Davies point.
@@
and The normalized vector is represented in FIG.(5) for $\phi=0.1$.
@@
The critical points are calculated by setting $\theta=\pi/2$ in $n^{1}$ and equating it to zero. We find two real and positive critical (or zero) points located at $(r_{+},\theta)=(0.57445,\pi/2),(0.99498,\pi/2)$ which are respectively named as $Z_{D}$ and $Z_{H}$. These points exactly match with the Davies points and the Hawking-Page phase transition point. We calculate the topological charge associated with each critical point by encircling contours around them as shown in FIG.(6). For the critical point $Z_{D}$ (which corresponds to the Davies point) the topological charge is $-1$ (FIG.(6(a)))and for the critical point $Z_{H}$ (which corresponds to the Hawking-Page phase transition point) the topological charge is $+1$(FIG.(6(a))).
Figure 6: RN AdS black hole : FIG. (a) represents topological number calculation for Davies point around the black contour in FiG.(5) and FIG. (b) represents the topological number calculation for Hawking-Page phase transition point around the blue contour in FIG.(5). We have considered $\phi=0.1$.
@@ V Kerr Ads Black hole
The metric for Kerr-Ads black hole in Boyer-Lindquist coordinate is given by :
$ds^{2}=-\frac{\Delta}{\rho^{2}}\left[dt-\frac{asin^{2}\theta d\varphi}{\Xi}\right]^{2}+\frac{\rho^{2}}{\Delta}dr^{2}+\frac{\rho^{2}}{\Sigma}d\theta^{2}+\frac{\Xi sin^{2}\theta}{\rho^{2}}\left[adt-\frac{r^{2}+a^{2}}{\Xi}d\varphi\right]^{2}$ (31)
where
$\Delta=(r^{2}+a^{2})\left(1+\frac{r^{2}}{l^{2}}\right)-2mr$
$\Sigma=1-\frac{a^{2}}{l^{2}}cos^{2}\theta\hskip 14.22636pt,\hskip 14.22636pt\rho^{2}=r^{2}++a^{2}cos^{2}\theta\hskip 14.22636pt,\hskip 14.22636pt\Xi=1-\frac{a^{2}}{l^{2}}$
Here $m$ is the mass parameter, $a$ is the angular momentum per unit mass and $l$ is the AdS length. The expressions for ADM mass $M$ and the angular momentum $J$ are :
$M=\frac{m}{\Xi^{2}}\hskip 14.22636pt,\hskip 14.22636ptJ=\frac{am}{\Xi^{2}}$ (32)
Solving $\Delta_{r}(r=r_{+})=0$ we can write mass parameter $m$ as :
$m=\frac{\left(a^{2}+r_{+}^{2}\right)\left(\frac{r_{+}^{2}}{l^{2}}+1\right)}{2r_{+}}$ (33)
The entropy of the black hole is given by :
$S=\frac{\pi(r_{+}^{2}+a^{2})}{\Xi}$ (34)
Using equation eq.(32) and eq.(34), the mass of Kerr Ads black hole in terms of entropy $S$ can be written as :
@@
Figure 7: Kerr AdS black hole: FIG. (a) represents $F$ vs $S$ plot for $\Omega=0.5$. The blue dot represents $S=3.62759$, the free energy becomes zero. This is the exact point at which the Hawking-Page transition takes place. FIG.(b) represents $C$ vs $S$ plot for $\Omega=0.5$. The black dot represents at $S=1.20764$, the specific heat diverges. This is the exact point at which the Davies-type transition takes place. In FIG.(c) spinodal curve is plotted for different values of $\Omega$.
@@
Here also the topological charge corresponding to the Davies point is determined to be $-1$, while for the Hawking-Page phase transition, it is $1$, as illustrated by the black and blue solid line respectively in FIG.(9).
@@ VI Conclusion
VII Acknowledgments
BH would like to thank DST-INSPIRE, Ministry of Science and Technology fellowship program, Govt. of India for awarding the DST/INSPIRE Fellowship[IF220255] for financial support.